find measurement of angle A(6x+7) and m of angle B(3x+20) if angle A and angle B are complementary
Answer by jim_thompson5910(35256)

angle A and angle B are complementary

m angle A + m angle B = 90

6x+7 + 3x+20 = 90

9x+27=90

9x=90-27

9x=63

x=(63)/(9)

x=7

Angle A: 6x+7 = 6(7)+7 = 49 degrees

Angle B: 3x+20 = 3(7)+20 = 41 degrees